OpenStudy (anonymous):

The sum of three consecutive odd intergers is 177. What are the intergers?

OpenStudy (anonymous):

n + n + 1 + n + 2 = 177. solve for n BUDDY

hartnn (hartnn):

nopes, n is odd so n+n+2+n+4 = 177

hartnn (hartnn):

@C$ can u find n ?

hartnn (hartnn):

3n+6=177 3n=177-6=171 n=171/3=57 so 3 odd numbers are 57,59,61